GraceWoods:

An advertisement claims that a 5-pound box of detergent will wash 20 loads of laundry. A 5-pound box costs $3.50. A 15-pound box costs $10.00. About how much will you save per load if you buy the 15-pound box?

KingKobra:

What are your answer choices?

KingKobra:

@Narad

Narad:

We must calculate the cost of one load if you buy the 5-pound box The cost is \[c _{1}=3.5/20=0.175$ /load\] If you buy the 15-pound box, the cost of one load is \[c _{2}=10/60=0.167 $/load\] The saving is \[=c _{1}-c _{2}=0.175-0.167=0.008$/load\]

GraceWoods:

Thanks guys